Uranium-235 has a half life of 7.04×10^8 years. How many years will it take for 99.9% of a U-235 sample to decay?
1
Expert's answer
2019-09-17T07:42:10-0400

1)     N(t) = N0 * 2 –t/T1/2

2)     0.1 = 100 * 2 –t/7.04*10^8

3)     t = 7.016 * 109